jagr2713 (jagr2713):

Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function

jagr2713 (jagr2713):

here is my function f(x)=6x+2 f(4)=f(4)+2=24+2=26

jagr2713 (jagr2713):

@zepdrix

jagr2713 (jagr2713):

@phi

OpenStudy (phi):

the first line, defining the function is ok the second line should just replace x with 4 (everywhere that you see x) so f(x)= 6x+2 evaluated at x=4 is written f(4) = 6*4 + 2 f(4)= 24+2 = 26

OpenStudy (phi):

your function is the equation of a line. A line has exactly one y value for every x value which means it is a function.

jagr2713 (jagr2713):

ok but now it ask me Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function do i prove something

OpenStudy (phi):

A function is a relation between (x,y) pairs where every x has exactly *one* y value associated with it.

OpenStudy (phi):

your function is the equation of a line. A line has exactly one y value for every x value which means it is a function.

OpenStudy (phi):

You could say, the function f(x)= 6x+2 has the form y= mx+b which is the equation of a line. A line has only 1 y value for each x value, so this equation represents a function.

jagr2713 (jagr2713):

oh ok

jagr2713 (jagr2713):

for this one Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences do u

jagr2713 (jagr2713):

take my regular function i used and replace 4 with 3 and explain the steps

OpenStudy (phi):

I would say, replace the x with 3

jagr2713 (jagr2713):

yea so it would be f(x)= 6x+2 f(3) = 6*3 + 2 f(3)= 20+2 = 22

OpenStudy (phi):

I guess your calculator is broken?

jagr2713 (jagr2713):

how

OpenStudy (phi):

3*6 is not 20

jagr2713 (jagr2713):

ohoho i added 2 already and added it again

jagr2713 (jagr2713):

my bad i forgot it there for the next line

OpenStudy (phi):

ok, just as long as you figured it out. So the answer is 20

jagr2713 (jagr2713):

yea sow i just explain the steps

jagr2713 (jagr2713):

and can i ask u 2 more questions please

OpenStudy (phi):

?

jagr2713 (jagr2713):

Using complete sentences, describe to the Martians how to find the inverse of your function.

OpenStudy (phi):

write the function as y= 6x+2 now "solve for x" can you do that ?

jagr2713 (jagr2713):

do you replace x with y?

OpenStudy (phi):

You can swap x and y, and then solve for y or you solve for x, then swap x and y

jagr2713 (jagr2713):

i really dont know how to do this inverse thing can you help

jagr2713 (jagr2713):

with an examlpe or something

OpenStudy (phi):

y= 6x+2 now "solve for x" start by adding -2 to both sides (that means write -2 on both sides of the equation) can you do that ?

jagr2713 (jagr2713):

y=x-2

jagr2713 (jagr2713):

y=1/3(x-2)

OpenStudy (phi):

y= 6x+2 write -2 on both sides

jagr2713 (jagr2713):

idk how to do this

jagr2713 (jagr2713):

y=6x-2

jagr2713 (jagr2713):

?

OpenStudy (phi):

almost. but you have an equation. You leave everything as it is but write -2 on the right side and -2 on the left side start with y= 6x+2 now write -2 on both sides: -2 + y = 6x +2 - 2

OpenStudy (phi):

the idea is we are "adding -2" or subtracting 2 from each side now simplify the right side. 2 -2 is ?

jagr2713 (jagr2713):

0

OpenStudy (phi):

and 6x+0 is 6x

jagr2713 (jagr2713):

-2+1 right

jagr2713 (jagr2713):

y=1

OpenStudy (phi):

-2 + y = 6x +2 - 2 becomes -2 + y = 6x +0 and that simplifies to -2 +y = 6x order does not matter when adding so we can also write y + -2 = 6x

OpenStudy (phi):

the x's and y's will not disappear. All we are doing is "moving" things around. so far we have 6x = y-2 now divide both sides by 6

OpenStudy (phi):

that means draw a line under the left side and write a 6 draw a line under the right side and write a 6

jagr2713 (jagr2713):

6x/6=y-2/6

OpenStudy (phi):

yes . but you should write the right side as (y-2)/6 because you are dividing (y-2) by 6 and the way you wrote it looks like y - (2/6) on the left side you have 6/6 . what is that ?

jagr2713 (jagr2713):

1

OpenStudy (phi):

so what do we have so far?

jagr2713 (jagr2713):

x=(y-2)/6

OpenStudy (phi):

we have "solved for x" (x is alone on one side) last step to find the inverse is swap the x and y

jagr2713 (jagr2713):

y=(x-2)/6

OpenStudy (phi):

yes

jagr2713 (jagr2713):

so that's the answer?

OpenStudy (phi):

yes, that is the inverse function

jagr2713 (jagr2713):

ty and my very last question

jagr2713 (jagr2713):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (phi):

you need to make up a new function

jagr2713 (jagr2713):

f(x)=4x+3

jagr2713 (jagr2713):

h(x)

OpenStudy (phi):

Then assign any number to x I would pick x=0 (keeps the arithmetic easy)

jagr2713 (jagr2713):

true h(0)

OpenStudy (phi):

now figure out f(h(x)) when x = 0 that means figure out h(0) and use that number for x in f(x)

jagr2713 (jagr2713):

h(x)=4x+3 h(0)=4*0+3 h(0)=4+3=7 f(x)= 6x+2 f(0) = 6*0 + 2 f(0)= 6+2 = 8

jagr2713 (jagr2713):

am i right

OpenStudy (phi):

you have some of the idea. Let's concentrate on h(0) once again your calculator is wrong. How did you get from the 2nd the the 3rd line for h(0) ?

jagr2713 (jagr2713):

i multiply 4*0=4+3=7

OpenStudy (phi):

4*0 = ?

jagr2713 (jagr2713):

000

jagr2713 (jagr2713):

i was rushing

jagr2713 (jagr2713):

h(x)=4x+3 h(0)=4*0+3 h(0)=0+3=3

OpenStudy (phi):

so h(0) = 3 now you want to find f( h(0) ) but h(0) is 3, so we want to find f(3)

jagr2713 (jagr2713):

f(x)= 6x+2 f(3) = 6*3 + 2 f(3)= 18+2 = 20

OpenStudy (phi):

that means f(h(0))= 20 now we need to find h(f(0)) first find f(0)

jagr2713 (jagr2713):

f(x)= 6x+2 f(0) = 6*0 + 2 f(0)= 6+2 = 8

OpenStudy (phi):

Gack!

jagr2713 (jagr2713):

huh

OpenStudy (phi):

redo what you posted and think about each step.

jagr2713 (jagr2713):

SAME MISTAKE

jagr2713 (jagr2713):

f(x)= 6x+2 f(0) = 6*0 + 2 f(0)= 0+2 = 2

OpenStudy (phi):

ok , f(0)= 2 can you find h(f(0)) ?

jagr2713 (jagr2713):

no

jagr2713 (jagr2713):

idk how to do that

OpenStudy (phi):

remember that f(0) is 2 replace f(0) with 2 in h(f(0)) what do you get ?

jagr2713 (jagr2713):

f(x)= 6x+2 f(2) = 6*2 + 2 f(2)= 12+2 = 14

OpenStudy (phi):

don't take so many steps. replace f(0) with 2 in h(f(0)) that means erase f(0) and put in a 2 write down what you get (don't do anything else, no math)

jagr2713 (jagr2713):

2=6x+2

OpenStudy (phi):

I want you to get the idea that h(f(0)) is h(2)

jagr2713 (jagr2713):

h(2)=6x+2 math is confusing

OpenStudy (phi):

yes, but if you understand each step, and go slow, you can learn it. remember, h(x) = 4x+3 , different from f(x)= 6x+2 notice h(2) means replace x with 2 in h(x)= 4x+3

jagr2713 (jagr2713):

h(x)=4x+3

jagr2713 (jagr2713):

h(2)=4x+3

OpenStudy (phi):

the x should have be replaced with a 2

jagr2713 (jagr2713):

i did that

OpenStudy (phi):

all x's h(2)=4x+3 <--- there is an x that should be a 2

jagr2713 (jagr2713):

h(2)=4(2)+3

OpenStudy (phi):

now simplify. 4x with x replaced by 2 is 4*2 4(2) means the same thing, but it's clearer to me to write it as 4*2 h(2)= 4*2 + 3 now simplify

jagr2713 (jagr2713):

h(2)=8+3=11